# Ontology Neural Network and ORTSF: A Framework for Topological Reasoning and Delay-Robust Control

### v2

@@ III Mathematical Preliminaries
This section introduces the mathematical concepts and tools underpinning the proposed Ontology Neural Network (ONN) and Ontological Real-Time Semantic Fabric (ORTSF). We provide precise definitions and interpretations of graph-based representations, discrete curvature, topological data analysis, delay-aware control dynamics, semantic map fusion, and ontology rule integration. These foundations are critical to the formal guarantees and design principles elaborated in later sections.
III-A Semantic Map and Graph Representations
A semantic map is defined as:
$\mathcal{M}=\{(p_{i},c_{i})\mid p_{i}\in\mathbb{R}^{3},c_{i}\in C\}$ (1)
The corresponding semantic graph is generated by:
$\mathcal{G}=\mathcal{F}_{\operatorname{graph}}(\mathcal{M})=(V,E,R)$ (2)
where:
$V$ $=\big{\{}v_{i}=(p_{i},c_{i})\big{\}},$
$E$ $=\big{\{}(v_{i},r_{ij},v_{j})\;\big{|}\;\operatorname{Cond}(v_{i},v_{j})\big{\}},$
$R$ $:E\to\mathcal{R}$ (3)
An edge is:
$e_{ij}=(v_{i},r_{ij},v_{j}),\quad r_{ij}\in\mathcal{R}$ (4)
III-B Forman-Ricci Curvature
$\operatorname{Ric}_{F}(e_{ij})=w(e_{ij})\bigg{[}$ $\frac{w(v_{i})+w(v_{j})}{w(e_{ij})}-\sum_{e_{k}\sim v_{i}}\frac{w(v_{i})}{\sqrt{w(e_{ij})w(e_{k})}}$
$-\sum_{e_{l}\sim v_{j}}\frac{w(v_{j})}{\sqrt{w(e_{ij})w(e_{l})}}\bigg{]}$ (5)
III-C Persistent Homology
$G_{t}^{0}\subseteq G_{t}^{\alpha_{1}}\subseteq\cdots\subseteq G_{t}^{\alpha_{n}}=G_{t}$ (6)
$d_{\operatorname{PH}}(D_{t},D_{t+\delta})=\inf_{\gamma}\sup_{x\in D_{t}}\|x-\gamma(x)\|_{\infty}$ (7)
@@ III-D Semantic Map Fusion
$T^{*}=\arg\min_{T}\sum_{(i,j)\in\mathcal{C}}\bigg{[}$ $\|T(p_{i}^{A})-p_{j}^{B}\|^{2}+\lambda\mathcal{L}(c_{i}^{A},c_{j}^{B})\bigg{]}$ (10)
@@ III-H Notation Summary
  • (1): semantic map
  • (2): map-to-graph transformation
  • (5): Forman-Ricci curvature
  • (7): PH distance
  • (57): delay model
  • (115): phase margin
  • (10): fusion objective
  • (18): ontology rule for planning
These formulations provide a rigorous basis for ONN reasoning, semantic consistency, and delay-compensated control in ORTSF.
IV Ontology Neural Network (ONN) Formalization
The Ontology Neural Network (ONN) constitutes the core reasoning architecture of the proposed framework, designed to encode, preserve, and reason over dynamic relational semantics in robotic perception and control. This section formalizes the ONN’s structure, its semantic state encoding, relational operators, loss formulation, and associated theoretical guarantees. Our formulation integrates geometric, topological, and functional semantics within a unified mathematical model.
IV-A Semantic State Tensor
Let $o_{i}$ denote an object in the scene at time $t$. Its semantic state is encoded by:
where:
  • $\mathbb{L}_{i}(t)$ represents locativeness: spatial position in a global or local reference frame.
  • $\mathbb{B}_{i}(t)$ represents boundedness: physical extent or affordance boundary.
  • $\mathbb{F}_{i}(t)$ represents formness: shape, appearance, or visual signature.
  • $\mathbb{I}_{i}(t)$ represents intentionality: inferred function, role, or task association.
These tensors form the feature vectors over which ONN operates, enabling reasoning that integrates physical, perceptual, and functional attributes.
IV-B Relational Encoding and Interaction Function
For object pair $(o_{i},o_{j})$, the relational descriptor is:
$R_{ij}(t)=\begin{bmatrix}d_{ij}(t)\\ \theta_{ij}(t)\\ \phi_{ij}(t)\end{bmatrix}$ (20)
where $d_{ij}$ is the Euclidean distance, $\theta_{ij},\phi_{ij}$ are orientation angles.
The interaction embedding is computed as:
$I_{ij}(t)=\mathcal{G}\big{(}\mathcal{S}_{i}(t),\mathcal{S}_{j}(t),R_{ij}(t)\big{)}$ (21)
where $\mathcal{G}$ may be instantiated as:
$\mathcal{G}=\begin{cases}\operatorname{MLP}([\mathcal{S}_{i},\mathcal{S}_{j},R_{ij}])\\ \operatorname{GCN}([\mathcal{S}_{i},\mathcal{S}_{j}],A)\\ \operatorname{Attention}(\mathcal{S}_{i},\mathcal{S}_{j},R_{ij})\end{cases}$ (22)
IV-C Scene Graph with Curvature Regularization
The scene is represented as:
$G_{\mathcal{C}}(t)=(V(t),E(t)),\quad V(t)=\{\mathcal{S}_{i}\},E(t)=\{I_{ij}\}$ (23)
Edges $E(t)$ are regularized by Forman-Ricci curvature:
$\operatorname{Ric}_{F}(e_{ij})=w(e_{ij})\bigg{[}$ $\frac{w(v_{i})+w(v_{j})}{w(e_{ij})}$
$-\sum_{e_{k}\sim v_{i}}\frac{w(v_{i})}{\sqrt{w(e_{ij})w(e_{k})}}$
$-\sum_{e_{l}\sim v_{j}}\frac{w(v_{j})}{\sqrt{w(e_{ij})w(e_{l})}}\bigg{]}$ (24)
IV-D Composite Loss Formulation
The ONN optimizes:
$\mathcal{L}_{\operatorname{total}}=\mathcal{L}_{\operatorname{pred}}+\lambda_{1}\mathcal{L}_{\operatorname{flow}}+\lambda_{2}\mathcal{L}_{\operatorname{relation}}+\lambda_{3}\mathcal{L}_{\operatorname{intent}}+\lambda_{4}\mathcal{L}_{\operatorname{context}}$ (25)
IV-D1 Multi-Objective Optimization Framework
The weighted sum approach in (25) can be extended to handle conflicting objectives more systematically using multi-objective optimization techniques.
$\varepsilon$-Constraint Formulation:
$\min_{\theta}\quad$ $\mathcal{L}_{\operatorname{pred}}$
subject to $\mathcal{L}_{\operatorname{context}}\leq\varepsilon_{c}$
$\mathcal{L}_{\operatorname{intent}}\leq\varepsilon_{i}$ (26)
Augmented Tchebycheff Method:
$\min_{\theta}\quad\max_{i}w_{i}(L_{i}-z_{i}^{*})+\rho\sum_{i}w_{i}(L_{i}-z_{i}^{*})$ (27)
where $z_{i}^{*}$ are ideal points and $\rho>0$ is a small augmentation parameter.
Pareto Optimality Analysis: The Karush-Kuhn-Tucker (KKT) conditions provide necessary optimality conditions:
$\nabla_{\theta}\mathcal{L}_{\operatorname{pred}}+\sum_{i}\mu_{i}\nabla_{\theta}g_{i}(\theta)$ $=0$ (28)
$\mu_{i}g_{i}(\theta)$ $=0,\quad\mu_{i}\geq 0$ (29)
where $g_{i}(\theta)=\mathcal{L}_{i}-\varepsilon_{i}$ are constraint functions.
The Lagrange multiplier path $\{\mu_{i}(\tau)\}$ traces the topology-control tradeoff surface as constraints $\{\varepsilon_{i}(\tau)\}$ vary.
Where:
$\mathcal{L}_{\operatorname{pred}}=\sum_{i}\|\hat{\mathcal{S}}_{i}(t+1)-\mathcal{S}_{i}(t+1)\|^{2}$ (30)
$\mathcal{L}_{\operatorname{flow}}=1-\frac{\dot{\mathcal{S}}_{i}(t)\cdot\dot{\hat{\mathcal{S}}}_{i}(t)}{\|\dot{\mathcal{S}}_{i}(t)\|\|\dot{\hat{\mathcal{S}}}_{i}(t)\|}$ (31)
$\mathcal{L}_{\operatorname{relation}}=\sum_{i,j}\|I_{ij}^{\operatorname{pred}}-I_{ij}^{\operatorname{GT}}\|^{2}$ (32)
$\mathcal{L}_{\operatorname{intent}}=-\sum_{c}y_{c}\log\hat{y}_{c}$ (33)
$\mathcal{L}_{\operatorname{context}}=\mathcal{L}_{\operatorname{ricci\text{-}internal}}+\lambda_{\operatorname{boundary}}\mathcal{L}_{\operatorname{ricci\text{-}boundary}}+\lambda_{\operatorname{ph}}\mathcal{L}_{\operatorname{ph}}$ (34)
IV-E Multi-Dimensional Topological Stability
IV-E1 High-Dimensional Persistent Homology Framework
For comprehensive topological analysis, we extend beyond 1-dimensional cycles to capture higher-dimensional voids and cavities. For each time $t$, we define a filtration function on edges:
$f_{t}:E(t)\to\mathbb{R},\quad f_{t}(e_{ij})=\alpha\|\mathcal{S}_{i}(t)-\mathcal{S}_{j}(t)\|_{2}+\beta|\operatorname{Ric}_{F}(e_{ij})|$ (35)
where $\alpha,\beta>0$ are weighting parameters.
We compute persistence diagrams $D_{k}(f_{t})$ for homological dimensions $k\in\{0,1,2,3\}$:
  • $H_{0}$: Connected components (object clustering)
  • $H_{1}$: 1-dimensional cycles (relational loops)
  • $H_{2}$: 2-dimensional cavities (enclosed regions, containers)
  • $H_{3}$: 3-dimensional voids (interior spaces, tunnels)
The multi-dimensional persistent homology distance is defined as:
$d_{\operatorname{PH}}^{(0:3)}(G_{\mathcal{C}}(t),G_{\mathcal{C}}(t+\delta))=\sum_{k=0}^{3}\alpha_{k}\,d_{B}\big{(}D_{k}(f_{t}),D_{k}(f_{t+\delta})\big{)}$ (36)
where $d_{B}$ is the bottleneck distance and $\{\alpha_{k}\}$ are dimension-specific weights with $\sum\alpha_{k}=1$.
IV-E2 Multi-Scale Filtration Framework
To capture topological features at multiple resolutions, we introduce a scale-space approach. Let $\Sigma=\{\sigma_{1}<\sigma_{2}<\cdots<\sigma_{m}\}$ be a set of scale parameters. For each scale $\sigma$, we define a smoothed filtration:
$f_{t}^{(\sigma)}(e_{ij})=\Phi_{\sigma}\big{(}f_{t}(e_{ij})\big{)}$ (37)
where $\Phi_{\sigma}$ is a smoothing operator (e.g., Gaussian convolution, morphological operations).
The multi-scale context loss becomes:
$\mathcal{L}_{\operatorname{context}}^{\operatorname{MS}}=\frac{1}{m}\sum_{\sigma\in\Sigma}\left(\mathcal{L}_{\operatorname{ricci}}^{(\sigma)}+\lambda_{\operatorname{ph}}\mathcal{L}_{\operatorname{ph}}^{(\sigma)}\right)$ (38)
This captures both fine-grained local relationships and coarse-grained global structure simultaneously.
IV-E3 Probabilistic Topological Stability
Under sensor noise and estimation uncertainties, we provide probabilistic guarantees for topological stability. Let $\xi_{t}$ represent the stochastic perturbation in the filtration function due to noise.
Assumption: The filtration perturbation $\|f_{t}-f_{t+\delta}\|_{\infty}$ is sub-Gaussian with parameter $\sigma^{2}$:
$\mathbb{E}\left[\exp\left(\frac{t(\|f_{t}-f_{t+\delta}\|_{\infty}-\mathbb{E}[\|f_{t}-f_{t+\delta}\|_{\infty}])}{\sigma}\right)\right]\leq\exp\left(\frac{t^{2}}{2}\right)$ (39)
Probabilistic Stability Theorem: Under the sub-Gaussian assumption, the persistent homology distance satisfies:
This provides confidence intervals for topological stability guarantees under realistic noise conditions.
Let $\overline{\operatorname{Ric}}_{F}$ be the mean curvature. We define:
$\mathcal{L}_{\operatorname{ph}}=\sum_{i,j}\operatorname{CE}(\hat{y}_{ij}^{\operatorname{sem}},y_{ij}^{\operatorname{sem}})$ (42)
where $\hat{y}_{ij}^{\operatorname{sem}}$ are predicted semantic labels and $y_{ij}^{\operatorname{sem}}$ are ground truth semantic labels for edge relationships. This formulation avoids circularity by separating topological distance from semantic label consistency.
Proposition: If $\mathcal{L}_{\operatorname{ricci\text{-}internal}}\to 0$ and $\mathcal{L}_{\operatorname{ph}}\to 0$, then the scene graphs exhibit topological stability:
$d_{\operatorname{PH}}(G_{\mathcal{C}}(t),G_{\mathcal{C}}(t^{\prime}))\to 0$ (43)
Note: This ensures filtration-level stability but does not guarantee graph isomorphism, which requires additional assumptions (same vertex sets, bijective mappings, homogeneous filtrations).
Proof (Sketch):
$\operatorname{Ric}_{F}(e)\approx\overline{\operatorname{Ric}}_{F}\Rightarrow E(t)\text{ has balanced curvature distribution}$ (44)
$\mathcal{L}_{\operatorname{ph}}\to 0\Rightarrow\text{semantic labels converge, reducing filtration perturbations}$ (45)
IV-F Theoretical Guarantee: Multi-Dimensional PH Stability Bound
Theorem (Multi-Dimensional PH Stability): Under the filtration function $f_{t}$ defined in (35), the multi-dimensional persistent homology distance satisfies:
$d_{\operatorname{PH}}^{(0:3)}\!\big{(}G_{\mathcal{C}}(t),G_{\mathcal{C}}(t+\delta)\big{)}$ $=\sum_{k=0}^{3}\alpha_{k}\,d_{B}\!\big{(}D_{k}(f_{t}),D_{k}(f_{t+\delta})\big{)}$ (46)
$\leq\sum_{k=0}^{3}\alpha_{k}\,\|f_{t}-f_{t+\delta}\|_{\infty}$ (47)
$\leq\sum_{k=0}^{3}\alpha_{k}\Big{(}C_{1,k}\,\kappa\,\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}\Big{)}$ (48)
$\quad+\sum_{k=0}^{3}\alpha_{k}\,C_{2,k}\,\eta(\mathcal{L}_{\operatorname{ph}}).$ (49)
where:
  • $d_{B}$ is the bottleneck distance between persistence diagrams $D(f_{t}),D(f_{t+\delta})$
  • $L_{c}>0$ is the Lipschitz constant relating curvature to filtration changes
  • $\kappa=\sqrt{|E|}\geq 1$ converts $L_{2}$ to $L_{\infty}$ norm (valid for finite graphs)
  • $\eta(\cdot)\geq 0$ accounts for semantic label mismatch effects on topology
This provides a formal bound linking ONN loss terms to topological stability.
IV-G Topological Neck Surgery Algorithm
To maintain topological stability and prevent degenerate connectivity, we introduce a formal neck surgery procedure that removes unstable topological features and replaces them with canonical structures.
Algorithm 1: Discrete Topological Neck Surgery
Input: Scene graph $G_{\mathcal{C}}(t)=(V,E)$, threshold $\epsilon_{\operatorname{neck}}>0$
Output: Surgically corrected graph $G^{\prime}_{\mathcal{C}}(t)$
Step 1: Neck Detection
for each 1-cycle $\gamma$ in $G_{\mathcal{C}}(t)$ do
  Compute persistence $\operatorname{pers}(\gamma)=\operatorname{death}(\gamma)-\operatorname{birth}(\gamma)$
  if $\operatorname{pers}(\gamma)<\epsilon_{\operatorname{neck}}$ then
   Mark $\gamma$ as unstable neck
  end if
end for
Step 2: Curvature-Based Validation
for each unstable neck $\gamma$ do
  $E_{\gamma}\leftarrow$ edges forming the neck cycle
  Compute $\operatorname{Ric}_{\operatorname{avg}}=\frac{1}{|E_{\gamma}|}\sum_{e\in E_{\gamma}}|\operatorname{Ric}_{F}(e)|$
  if $\operatorname{Ric}_{\operatorname{avg}}>\overline{\operatorname{Ric}}_{F}+2\sigma_{\operatorname{Ric}}$ then
   Confirm $\gamma$ for surgery (high curvature indicates geometric instability)
  end if
end for
Step 3: Canonical Replacement
for each confirmed neck $\gamma$ do
  $V_{\gamma}\leftarrow$ vertices involved in neck cycle
  Remove edges $E_{\gamma}$ from $G_{\mathcal{C}}(t)$
  if $|V_{\gamma}|=3$ then
   Insert triangle clique: $E_{\operatorname{new}}=\{(v_{i},v_{j})|v_{i},v_{j}\in V_{\gamma},i\neq j\}$
  else if $|V_{\gamma}|=4$ then
   Insert star graph: Select central vertex $v_{c}\in V_{\gamma}$,
    $E_{\operatorname{new}}=\{(v_{c},v)|v\in V_{\gamma}\setminus\{v_{c}\}\}$
  else
   Insert minimum spanning tree on $V_{\gamma}$
  end if
  Update edge weights: $w(e_{\operatorname{new}})=\operatorname{median}\{w(e)|e\in E_{\gamma}\}$
end for
Step 4: Monotonic Verification
 Compute $\mathcal{L}_{\operatorname{ricci\text{-}internal}}^{\operatorname{new}}$ and $\lambda_{2}^{\operatorname{new}}$ (Fiedler eigenvalue)
if $\mathcal{L}_{\operatorname{ricci\text{-}internal}}^{\operatorname{new}}\leq\mathcal{L}_{\operatorname{ricci\text{-}internal}}$ and $\lambda_{2}^{\operatorname{new}}\geq\lambda_{2}$ then
  Accept surgery: $G^{\prime}_{\mathcal{C}}(t)\leftarrow$ updated graph
else
  Reject surgery: $G^{\prime}_{\mathcal{C}}(t)\leftarrow G_{\mathcal{C}}(t)$ (original)
end if
Theoretical Guarantees:

### v3

@@ I Introduction
Figure 1: ONN + Contextual Topology Framework Overview. The complete architectural pipeline shows the flow from existence recognition through ONN processing, 1-cochain output, cycle constraint projection (CPL), connection-consistency resolution (CCL), to ORTSF delay compensation. Each component implements specific theoretical guarantees from the seven-theorem framework, providing mathematical rigor for semantic reasoning and control integration.
@@ III Mathematical Preliminaries
This section establishes the rigorous mathematical foundations for dynamic context preservation in robotic semantic reasoning. We formalize the central challenge of maintaining contextual topology under continuous environmental changes through a unified framework of graph cochains, contextual topology constraints, and averaged operator theory. These theoretical foundations underpin the seven core theorems that guarantee convergence, stability, and context preservation in the ONN+ORTSF architecture.
III-A Fundamental Definitions: Graphs and 1-Cochains
Definition III.1 (Graph and 1-Cochain Space).
Let $G=(V,E)$ be a connected undirected graph with $|V|=n$ vertices and $|E|=m$ edges. Fix an arbitrary orientation for edges to define the boundary operator $B_{1}\in\mathbb{R}^{m\times n}$. The edge-signal (1-cochain) space $\mathcal{X}=\mathbb{R}^{m}$ is endowed with the standard inner product, forming a Hilbert space structure.
Definition III.2 (Edge Laplacian and Harmonic Subspace).
The edge Laplacian is defined as:
$L_{1}:=B_{1}B_{1}^{\top}\succeq 0$ (1)
The harmonic (loop) subspace $\mathcal{H}:=\ker L_{1}$ satisfies $\mathcal{H}=\operatorname{Im}(B_{1})^{\perp}$ in the orthogonal decomposition $\mathcal{X}=\operatorname{Im}(B_{1})\oplus\operatorname{Im}(B_{1})^{\perp}$.
Definition III.3 (Contextual Topology: Cycle Constraints).
Given a spanning tree $T$ and fundamental cycles $\{\ell_{1},\ldots,\ell_{q}\}$ from non-tree edges, let $C\in\mathbb{R}^{q\times m}$ be the loop-edge signature matrix. The contextual topology constraint is:
$\mathcal{C}:=\{x\in\mathcal{X}\mid Cx=\tau\},\quad\tau\in\mathbb{R}^{q}$ (2)
This encodes ”loops/configurations that must not be broken” as linear equality constraints.
Definition III.4 (Connection Consistency: Connection Laplacian).
For node variables $f=(f_{i})_{i\in V}\in\mathbb{R}^{dn}$, edge transformations $T_{ij}\in\mathbb{R}^{d\times d}$, and edge weights $w_{ij}>0$, the global consistency minimization is:
$\min_{f}\Phi(f):=\frac{1}{2}\sum_{(i,j)\in E}w_{ij}\|T_{ij}f_{j}-f_{i}\|^{2}$ (3)
The normal equations have the form $L_{\text{conn}}f=b$, with unique solution under gauge anchoring.
III-B Dynamic Cochain Extensions
Definition III.5 (Piecewise-Static Cochain Bundle).
Partition the time axis into intervals $t_{k},t_{k+1})$ with fixed simplicial complexes $K_{k}$ on each interval. Define time-parametrized cochains $c_{t}:K_{t}^{(k)}\to\mathbb{R}$ for $t\in\mathbb{R}_{+}$. Transition maps $\phi_{k}:C^{k}(K_{k})\to C^{k}(K_{k+1})$ handle simplicial changes at interval boundaries.
Definition III.6 (Contextual Invariance Condition).
For all time $t$, the contextual topology constraint must be preserved:
$\Psi(t):\mathcal{X}(t)\to\mathbb{R}^{q},\quad\Psi(c_{t})=0\text{ for all }t$ (4)
where $\Psi$ is a smooth constraint functional ensuring specific relationships remain invariant despite structural changes.
III-C Forman-Ricci Curvature and Filtrations
The Forman-Ricci curvature for edge $e_{ij}\in E$ following Forman’s discrete Bochner method [7]:
$\operatorname{Ric}_{F}(e_{ij})$ $=w(e_{ij})\bigg{[}\frac{w(v_{i})+w(v_{j})}{w(e_{ij})}-\sum_{e_{k}\sim v_{i},e_{k}\neq e_{ij}}\frac{w(v_{i})}{\sqrt{w(e_{ij})w(e_{k})}}$
$\qquad-\sum_{e_{l}\sim v_{j},e_{l}\neq e_{ij}}\frac{w(v_{j})}{\sqrt{w(e_{ij})w(e_{l})}}\bigg{]}$ (5)
where $w(v_{i})$ and $w(e_{ij})$ are positive weights assigned to vertices and edges. This discrete curvature measures the ”bending” of the edge within the combinatorial structure [8].
For temporal analysis, we construct a filtration based on the semantic-geometric function $f_{t}:E\to\mathbb{R}$:
$f_{t}(e_{ij})=\alpha\|\mathcal{S}_{i}(t)-\mathcal{S}_{j}(t)\|_{2}+\beta|\operatorname{Ric}_{F}(e_{ij})|$ (6)
The filtration is given by sublevel sets:
$G_{t}^{0}\subseteq G_{t}^{\alpha_{1}}\subseteq\cdots\subseteq G_{t}^{\alpha_{n}}=G_{t},\quad G_{t}^{\alpha_{k}}=\{e\in E:f_{t}(e)\leq\alpha_{k}\}$ (7)
where $0<\alpha_{1}<\alpha_{2}<\cdots<\alpha_{n}$ [9].
The bottleneck distance between persistence diagrams:
$d_{\operatorname{PH}}(D_{t},D_{t+\delta})=d_{B}(D_{t},D_{t+\delta})=\inf_{\gamma:D_{t}\to D_{t+\delta}}\sup_{x\in D_{t}}\|x-\gamma(x)\|_{\infty}$ (8)
where $\gamma$ ranges over all bijections from $D_{t}$ to $D_{t+\delta}$, extended to include the diagonal $\{(x,x):x\in\mathbb{R}\}$ [9].
@@ III-D Semantic Map Fusion
$T^{*}$ $=\arg\min_{T}\sum_{(i,j)\in\mathcal{C}}\bigg{[}\|T(p_{i}^{A})-p_{j}^{B}\|^{2}$
$\qquad+\lambda\mathcal{L}(c_{i}^{A},c_{j}^{B})\bigg{]}$ (11)
@@ III-H Notation Summary
Key mathematical constructs established in this section:
  • Graph structures: $G=(V,E)$ with boundary operator $B_{1}$ and edge Laplacian $L_{1}=B_{1}B_{1}^{T}$
  • Cochain spaces: Edge-signal space $\mathcal{X}=\mathbb{R}^{m}$ with Hilbert structure
  • Contextual constraints: $\mathcal{C}=\{x\in\mathcal{X}\mid Cx=\tau\}$ via (2)
  • Connection consistency: Minimization $\min_{f}\Phi(f)$ from (3)
  • Forman-Ricci curvature: Discrete curvature measure via (5)
  • Persistent homology: Topological distance $d_{\operatorname{PH}}$ from (8)
  • Delay dynamics: Plant model $G_{d}(s)=G(s)e^{-s\Delta t}$ from (68)
  • Semantic tensors: Node representations $\mathcal{S}_{i}(t)\in\mathbb{R}^{d}$ from (16)
  • Ontological rules: Logic constraints via (18) and (19)
These mathematical foundations provide the rigorous basis for the seven-theorem framework, ensuring that ONN reasoning, contextual topology preservation, and ORTSF delay compensation operate within well-defined functional analytic structures.
IV Main Theoretical Results
This section establishes the seven core theorems that provide rigorous mathematical guarantees for dynamic context preservation in the ONN+ORTSF framework. These theorems collectively ensure that contextual topology is preserved under temporal evolution, structural changes, and system delays, while maintaining convergence and stability properties essential for reliable robotic operation.
IV-A Core Mathematical Framework: Context-Preserving Operations
The ONN-interpretation-control pipeline processes ONN-generated 1-cochains $\hat{x}\in\mathcal{X}$ and local estimates $\hat{f}$ through two context-preserving layers: